LOG-IN
Displaying report 1-1 of 1.
MIF (General)
Print this report.
yutaro.enomoto - 22:47, Sunday 18 November 2018 (7030)Get code to link to this report
Some measurements on ALS CARM

= Crossover frequency measurement =
Mesured the TF from LSC-MCL_IN2 to LSC-MCL_IN1 by injecting sine signal to LSC-MCL_EXC.
The current setings for ALS CARM is, IN2 gain = 16dB, FAST gain = 12 dB, and LSC-MCL_GAIN is 1.2.

= Power spectrum of the error signal of ALS CARM =
The Y axis is roughly calibrated to Hz/sqrt(Hz), including the effect of cavity pole of the arm.
Suppression around several hundred Hz seems to be not sufficient???

= REFLAIR A RF45 I is fluctuating =
The "offset" level is fluctuating. Probably due to alignment fluctuation of some mirrors. Note that I have tried to optimize the alignment of main laser to X arm by looking at the transmission of 10/01 modes.
The amplitude of the fluctuation is larger than the amplitude of PDH signal, so the locking the arm with this signal is going to be impossible unless this fluctuation is removed.

Images attached to this report
7030_20181118142331_alscarmcross.png 7030_20181118142937_errorsignalalscarm.png 7030_20181118144725_rf45fluctuating.png